# bibfile.bib

### v1

@@ Abstract
Let $\mathbb{F}$ be a non-archimedean local field, and let $G=\mathrm{SL}_{n}(\mathbb{F})$, $n\geq 3$. Let $X$ be an infinite-dimensional simply connected symmetric space of finite rank, with non-positive curvature operator. We prove that every continuous action by isometries of $G$ on $X$ has a fixed point.
@@ 1 Introduction
The main goal of this paper is to study continuous actions by isometries of the group $G=\mathrm{SL}_{n}(\mathbb{F})$, with $\mathbb{F}$ a non-archimedean local field, on simply connected infinite-dimensional symmetric spaces $X$ of non-positive curvature and finite rank, where the rank is defined as the dimension of a maximal embedded Euclidean subspace. Such spaces come from Hilbert spaces with indefinite bilinear forms, whose study dates back to Pontrjagin [pontrjagin], Iohvidov-Krein [iohkrein], and Naimark [naimark1, naimark2, naimark3, naimark4]. The study of the associated symmetric spaces was first suggested by Gromov ([gromov], Section 6). The particular case of the infinite-dimensional real hyperbolic space has been treated by Burger, Iozzi and Monod in [bim], where they construct equivariant embedding of trees into this space. More recently, infinite-dimensional symmetric spaces of finite rank have been widely studied by Duchesne [ducphd, duc2013, duc2015, duc2023].
In another paper [trees], we have studied actions of tree automorphism groups on such spaces $X$, showing that they can be irreducible only if $X$ has rank one. Here, we focus on actions of higher-rank algebraic groups, such as $G=\mathrm{SL}_{n}(\mathbb{F})$ with $n\geq 3$, on the same spaces $X$.
Our main goal will be to prove that, if $n\geq 3$, the group $G$ does not have any continuous fixed point-free action by isometries on $X$.
This is in contrast to the case $n=2$, where there exist continuous fixed point-free actions of $\mathrm{SL}_{2}(\mathbb{F})$ by isometries on the infinite-dimensional real hyperbolic space ([bim], Theorem C).
We start by precisely defining the family $\mathcal{X}$ of possible target spaces.
@@
We have to include finite-dimensional symmetric spaces in the family $\mathcal{X}$, even though they are not the main focus of our paper, because we will prove our result by an induction that also relies on the finite-dimensional case.
@@
We can further say that, in the case where the space $X$ is finite-dimensional, the problem is already well understood. Indeed, let $f:G\rightarrow\text{Isom}(X)$ be any continuous homomorphism. Since $\text{Isom}(X)$ is a Lie group it has no small subgroups, and since $G$ is totally disconnected, by Van Dantzig’s theorem there is a basis of neighborhoods of the identity that are open compact subgroups. This implies that there exists an open compact subgroup of $G$ whose image in $\text{Isom}(X)$ is trivial. As $\mathrm{PSL}_{n}(\mathbb{F})$ is simple ([grove], Theorem 1.13), and the kernel of the projection of $\mathrm{SL}_{n}(\mathbb{F})$ on $\mathrm{PSL}_{n}(\mathbb{F})$ is finite, this implies that the homomorphism $f$ is trivial on a finite-index subgroup of $G$, i. e. the homomorphism has finite image. Then, since $X$ is a CAT(0) space, the existence of a bounded orbit implies the existence of a fixed point ([bh], II.2.8).
The first result in the infinite-dimensional case is due to Burger, Iozzi and Monod ([bim], Theorem C). In that paper the authors construct a family of equivariant embeddings of a regular tree $T_{d}$ (a tree where every vertex has degree $d\geq 3$) into an infinite-dimensional hyperbolic space, which corresponds to a family of continuous irreducible representations of the group $\mathrm{Aut}(T_{d})$ to $O(1,\infty)$, i. e. continuous irreducible representations to a separable Hilbert space that preserve a non-degenerate bilinear form of index $1$. This implies in particular that for every non-archimedean local field $\mathbb{F}$ the group $\mathrm{SL}_{2}(\mathbb{F})$, which is a subgroup of $\mathrm{Aut}(T_{q+1})$ where $q$ is the cardinality of the residue field of $\mathbb{F}$, has continuous fixed point-free actions by isometries on the infinite-dimensional real hyperbolic space.
In this paper, we focus on the higher-rank case, namely $G=\mathrm{SL}_{n}(\mathbb{F})$ with $n\geq 3$. We will show that the behavior seen in [bim] for $\mathrm{SL}_{2}(\mathbb{F})$ cannot happen in this case. Arturo Sánchez Gonzáles obtained a first result in this direction in his Ph.D. thesis, where he proved that for $n\geq 3$ there are no continuous irreducible representations of $\mathrm{SL}_{n}(\mathbb{Q}_{p})$ to $\mathrm{O}(2,\infty)$ ([arturo], Teorema C).
Our main result is the following:
Theorem 1.4.
Let $\mathbb{F}$ be a non-archimedean local field, and let $n\geq 3$. Let $G=\mathrm{SL}_{n}(\mathbb{F})$. Let $X\in\mathcal{X}$, as defined in Definition 1.1. Then any continuous action of $G$ by isometries on $X$ fixes a point in $X$.
Remark 1.5.
We can already observe that, if we can show the theorem for $n=3$, then the general case follows. In fact, it is shown in [cornulier] (Theorem 1.4, Proposition 1.2) that if $G$ is a simple linear algebraic group over a local field, then every action by isometries of $G$ on any metric space $X$ is either bounded, which means that the orbits are bounded, or metrically proper, which means that for every bounded subset $B\subset X$ the set $\{g\in G:gB\cap B\neq\emptyset\}$ has compact closure. We apply the result to $G=\mathrm{SL}_{n}(\mathbb{F})$. In case the action is bounded, we can already deduce that it has a fixed point in $X$, since our target spaces are CAT(0) ([bh], II.2.8). In case it is metrically proper, we take a closed subgroup $H<G$ isomorphic to $\mathrm{SL}_{3}(\mathbb{F})$. If $H$ has a fixed point $x\in X$, we know that the set $\{g\in G:gx=x\}$ has compact closure, which is a contradiction since it contains $H$ which is closed and non-compact.
Remark 1.6.
It is enough to show that the action of a cocompact subgroup $\Gamma<G$ is bounded. We will take $\Gamma$ to be a torsion-free cocompact lattice of $G$, that is, a torsion-free discrete cocompact subgroup of $G$. The existence of a discrete cocompact subgroup $\Gamma^{\prime}<G$ was proved in [borelharder] (Theorem A, Theorem 3.3), and due to ([garland], Theorem 2.7), there exists $\Gamma\lhd\Gamma^{\prime}$ of finite index that is also torsion-free.
Remark 1.7.
Remark 1.5 allows to extend Theorem 1.4 to a more general class of groups $G$, namely the class of all simple linear algebraic groups over a local field that contain $\mathrm{SL}_{3}(\mathbb{F})$ as a closed subgroup.
We conclude our introduction with a short outline of the paper.
After all the above reductions, our goal is to prove Theorem 1.4 for a torsion-free cocompact lattice $\Gamma$ in $G=\mathrm{SL}_{3}(\mathbb{F})$. First, we will show (Section 2) that every continuous action of $\Gamma$ by isometries on $X\in\mathcal{X}$ has a fixed point in $\overline{X}=X\cup\partial X$, where $\partial X$ is the boundary at infinity of $X$. This is done by generalizing a result that was proved by Wang [wang] in the finite-dimensional case. Finally, we will show (Section 3) that if a group $\Gamma$ with Kazhdan’s Property (T) has the property that whenever it acts continuously by isometries on a space $X\in\mathcal{X}$ it fixes a point in $\overline{X}$, and all its finite-index subgroups have the same property, then the group actually has to fix a point in $X$. This will allow us to conclude our proof.
@@
This paper is part of my Ph.D. project under the supervision of Prof. Nicolas Monod at EPFL (Lausanne, Switzerland). I am grateful to Nicolas for proposing this project, for giving me valuable advice and feedback, and for sharing with me deep insights of the theory. I am also grateful to Bruno Duchesne and Pierre Py for their useful comments and suggestions on this paper.
@@ 2 Fixing a point in $\overline{X}$
Let $\mathbb{F}$ be a non-archimedean local field, let $G=\mathrm{SL}_{3}(\mathbb{F})$, and let $\Gamma$ be a discrete, torsion-free, cocompact subgroup of $G$. Let $X\in\mathcal{X}$. Then every continuous action of $\Gamma$ by isometries on $X$ fixes a point in $\overline{X}$.
This result was proved by Wang in [wang] in the case where $X$ is a finite-dimensional Riemannian manifold of non-positive sectional curvature. The proof of our Theorem 2.1 will closely follow the proof in [wang], with some adaptations: most importantly, we need to take into account the fact that our space $X\in\mathcal{X}$ might also be infinite-dimensional.
@@ 2.1 The building of $G$ and the eigenvalue property
First, we introduce the Bruhat-Tits building $\Sigma$ of the group $G$, which is a Euclidean building of dimension two on which $G$ acts by isometries. It is a union of infinitely many Euclidean planes (apartments), each of them being tessellated by equilateral triangles (chambers). The action of $G$ is transitive in the set of pairs $(\mathcal{A},\mathcal{C})$ where $\mathcal{A}$ is an apartment and $\mathcal{C}$ is a chamber contained in $\mathcal{A}$. More details on the construction and structure of the building are explained in [buildings] (6.9).
@@
We describe the structure of the graph $Lk(v)$. As mentioned in [buildings] (6.9), it is the flag complex of $\mathrm{SL}_{3}(k)$, where $k$ is a finite field: the residue field of $\mathbb{F}$. This means that the vertices are points and lines in the projective plane over $k$, and whenever a point is contained in a line, the corresponding vertices are connected by an edge. If $q$ is the cardinality of the field $k$ (a power of a prime number), there are $N=2(q^{2}+q+1)$ vertices in total: $q^{2}+q+1$ points of $\mathbb{P}^{2}(k)$ and $q^{2}+q+1$ lines. Each vertex is connected to $q+1$ other vertices, thus there are in total $M=(q+1)(q^{2}+q+1)$ edges.
We show a property of the graph $Lk(v)$ that corresponds to the first eigenvalue property in [wang]. Let $v\in\Sigma(0)$ and let $Lk(v)$ be its link graph. After choosing an orientation for every edge in $Lk(v)(1)$, for every $f:Lk(v)(0)\rightarrow\mathbb{R}$ we can define $df:Lk(v)(1)\rightarrow\mathbb{R}$ by $df((u,w))=f(u)-f(w)$. We let $||f||^{2}=\sum_{u\in Lk(v)(0)}|f(u)|^{2}$ and $||df||^{2}=\sum_{(u,w)\in Lk(v)(1)}|f((u,w))|^{2}$.
Lemma 2.2.
There exists $\lambda>\frac{q+1}{2}$ such that, if $f$ satisfies $\sum_{u\in Lk(v)(0)}{f(u)}=0$, then $||df||^{2}\geq\lambda||f||^{2}$.
Proof.
Let $A$ be the adjacency matrix of $Lk(v)$. More explicitly, it is a matrix with $M$ rows and $N$ columns where, after an ordering of the vertices and of the edges has been chosen, the coefficient $A_{ij}$ is $1$ if the $i$-th edge starts from the $j$-th vertex, $-1$ if the $i$-th edge terminates at the $j$-th vertex, and 00 in all other cases. Then we can consider the function $f$ as a vector in $\mathbb{R}^{N}$ and the function $df$ as a vector in $\mathbb{R}^{M}$, so that we have $df=Af$. Therefore we have
$||df||^{2}=\langle df,df\rangle=\langle Af,Af\rangle=\langle f,A^{T}Af\rangle,$
@@
In order to bound $\langle A^{T}Af,f\rangle$ in terms of $\langle f,f\rangle=||f||^{2}$, we study the eigenvalues of the matrix $B=A^{T}A$. It is a $N\times N$ square matrix where the coefficient $B_{ij}$ is $q+1$ if $i=j$, $-1$ if $i\neq j$ and the $i$-th and $j$-th vertex are connected by an edge, 00 if $i\neq j$ and the $i$-th and $j$-th vertex are not connected. We now describe the eigenspaces of $B$ with their respective eigenvalues:
  • If $f$ is a constant vector, then $Bf=0$ and thus $f$ is an eigenvector of $B$ with eigenvalue 00. This makes a $1$-dimensional eigenspace.
  • If $f$ takes a constant value $c$ on all points $p$ of the projective plane $\mathbb{P}_{2}(k)$, and the value $-c$ on all lines $l$ in $\mathbb{P}_{2}(k)$, then (recall that the graph is bipartite with respect to points and lines) $Bf=2(q+1)f$ and thus $f$ is an eigenvector of $B$ with eigenvalue $2(q+1)$. This makes another $1$-dimensional eigenspace.
  • Let $f$ be defined arbitrarily on the set of points $p$, with sum zero on this set. Then define $f$ at each line $l$ as $f(l)=\frac{1}{\sqrt{q}}\sum_{p\in l}{f(p)}$. Observe that, for every point $p$, we have
    $\frac{1}{\sqrt{q}}\sum_{l\ni p}{f(l)}=\frac{1}{q}\sum_{l\ni p}{\sum_{p^{\prime}\in l}{f(p^{\prime})}}=\frac{1}{q}\left(qf(p)+\sum_{p^{\prime}\in\mathbb{P}_{2}(k)}{f(p^{\prime})}\right)=f(p),$
    since when we sum over all points in all lines passing through $p$ the point $p$ is summed $q+1$ times while all the other points are summed once. We can conclude that, for every vertex $u\in Lk(v)(0)$, the sum of $f$ at all neighbors of $u$ in the graph is equal to $\sqrt{q}f(u)$. This implies that $Bf=(q+1-\sqrt{q})f$. As $f$ was chosen arbitrarily with sum zero on the set of points $p\in\mathbb{P}_{2}(k)$, this makes a $(q^{2}+q)$-dimensional eigenspace with eigenvalue $q+1-\sqrt{q}$.
  • In the previous construction, after fixing $f$ arbitrarily with sum zero on the set of points $p$ in $\mathbb{P}_{2}(k)$, we might also choose to define it on the set of lines by setting $f(l)=-\frac{1}{\sqrt{q}}\sum_{p\in l}{f(p)}$ for every line $l$. In this case we obtain $f(p)=-\frac{1}{\sqrt{q}}\sum_{l\ni p}{f(l)}$ for every point $p$, hence a $(q^{2}+q)$-dimensional eigenspace with eigenvalue $q+1+\sqrt{q}$.
With the above description, we have decomposed $\mathbb{R}^{N}$ as orthogonal direct sum of four eigenspaces for the operator $B$. What we need to show for the lemma is that, whenever $f$ is a vector in $\mathbb{R}^{N}$ with sum of the coordinates equal to zero, we have $\langle Bf,f\rangle\geq\lambda||f||^{2}$ for some $\lambda>\frac{q+1}{2}$. We can pick $\lambda=q+1-\sqrt{q}$, which is greater than $\frac{q+1}{2}$ for $q\geq 2$. Since the sum of the coordinates is zero the vector $f$ decomposes as an orthogonal sum of vectors in the second, third and fourth eigenspace described above, whose eigenvalues are all greater than or equal to $\lambda$. Hence the result follows.
@@
The function $f$ in Lemma 2.2 takes values in $\mathbb{R}$, but the result also holds when it takes values in a general Hilbert space. To see this, it suffices to restrict to the finite-dimensional subspace spanned by the (finitely many) values of $f$, then apply the lemma component-wise.

### v2

# A fixed point theorem for the action of linear higher rank algebraic groups over local fields on symmetric spaces of infinite dimension and finite rank

@@ Abstract
Let $\mathbb{F}$ be a non-archimedean local field of characteristic zero whose residue field has at least three elements. Let $G$ be an almost simple linear algebraic group over $\mathbb{F}$, with $\mathrm{rank}_{\mathbb{F}}(G)\geq 2$. Let $X$ be a simply connected symmetric space of infinite dimension and finite rank, with non-positive curvature operator. We prove that every continuous action by isometries of $G$ on $X$ has a fixed point. If the group $G$ contains $\mathrm{SL}_{3}(\mathbb{F})$, the result holds without any assumption on the non-archimedean local field $\mathbb{F}$. The result extends to cocompact lattices in $G$ if the cardinality of the residue field of $\mathbb{F}$ is large enough, with a bound that depends on $\mathrm{rank}_{\mathbb{F}}(G)$.
@@ 1 Introduction
The theory of finite-dimensional or unitary representations of algebraic groups is a very classical subject. A topic of much recent interest is representations preserving a sesquilinear form of finite index on a Hilbert space, or representations into Pontryagin spaces [39]. These are particularly interesting because their projective versions appear as isometry groups of non-positively curved symmetric spaces of infinite dimension and finite rank ([19], Theorem 3.3).
In the case of Lie groups, which include linear algebraic groups over $\mathbb{R}$ or $\mathbb{C}$, there are examples of continuous irreducible representations of groups of rank one into Pontryagin spaces of arbitrary index [14, 33], and in contrast no such representation exists for groups of higher rank ([19], Theorem 1.1). In the case of linear algebraic groups over non-archimedean local fields, there exist continuous irreducible representations of groups of rank one, such as $\mathrm{SL}_{2}(\mathbb{Q}_{p})$, into the isometry group of the infinite-dimensional real hyperbolic space ([11], Theorem C), that is, the projective real Pontryagin space of index one.
In this paper, we focus on the case of linear higher rank algebraic groups over non-archimedean local fields. We will prove the following theorem:
Theorem 1.1.
Let $\mathbb{F}$ be a non-archimedean local field of characteristic zero whose residue field $k$ has at least three elements. Let $G$ be an almost simple linear algebraic group over $\mathbb{F}$, with $\mathrm{rank}_{\mathbb{F}}(G)\geq 2$. Let $X$ be an infinite-dimensional simply connected symmetric space with finite rank and non-positive curvature operator. Then, any continuous action of $G$ by isometries on $X$ fixes a point in $X$.
The rank of a symmetric space $X$ is defined as the maximal dimension of an isometrically embedded Euclidean space in $X$. The theorem implies that there are no continuous irreducible representations of $G$ into any Pontryagin space.
A first result in this direction was obtained by Arturo Sánchez Gonzáles in his Ph.D. thesis, where he proved that for $n\geq 3$ there are no continuous irreducible representations of $\mathrm{SL}_{n}(\mathbb{Q}_{p})$ to $\mathrm{O}(2,\infty)$ ([26], Teorema C).
It is known that any almost simple linear algebraic group $G$ over a non-archimedean local field $\mathbb{F}$ with $\mathrm{rank}_{\mathbb{F}}(G)\geq 2$ contains a quotient of $\mathrm{SL}_{3}(\mathbb{F})$ or $\mathrm{Sp}_{4}(\mathbb{F})$ by a finite subgroup (Proposition I.1.6.2 in [32], which is a consequence of Theorem 7.2 in [7]). The strategy for the proof of Theorem 1.1 will be to restrict ourselves to these two cases and to use a cocompact lattice $\Gamma$ in $G=\mathrm{SL}_{3}(\mathbb{F})$ or $G=\mathrm{Sp}_{4}(\mathbb{F})$ to build $\Gamma$-equivariant functions from the Bruhat-Tits building of $G$ to $X$. The existence of a cocompact lattice follows from ([6], Theorem A) since the field $\mathbb{F}$ has characteristic zero. This is why we need to make that assumption in the theorem: if the field $\mathbb{F}$ has positive characteristic, then a cocompact lattice exists in $\mathrm{SL}_{3}(\mathbb{F})$ ([6], Theorem 3.3) but not in $\mathrm{Sp}_{4}(\mathbb{F})$ ([32], Corollary IX.4.8), therefore, we cannot carry out our proof in that case. However, we are able to solve the case of $\mathrm{SL}_{3}(\mathbb{F})$ in any characteristic and, as we will see in the proof, this case also does not require the assumption on the residue field. This leads to the following result:
Theorem 1.2.
Let $\mathbb{F}$ be a non-archimedean local field. Let $G$ be an almost simple linear algebraic group over $\mathbb{F}$, containing a quotient of $\mathrm{SL}_{3}(\mathbb{F})$ by a finite subgroup. Let $X$ be an infinite-dimensional simply connected symmetric space with finite rank and non-positive curvature operator. Then, any continuous action of $G$ by isometries on $X$ fixes a point in $X$.
@@
The study of Hilbert spaces with indefinite sesquilinear forms dates back to Pontryagin [39], Iohvidov-Krein [30], and Naimark [35, 36, 37, 38]; further treatment can be found in [2, 5, 29]. The study of the associated symmetric spaces was first suggested by Gromov ([27], Section 6). The particular case of the infinite-dimensional real hyperbolic space $\mathbf{H}^{\infty}$ has been treated by Burger, Iozzi, and Monod in [11], where they construct equivariant embedding of trees into this space and classify continuous irreducible representations of groups such as $\mathrm{SL}_{2}(\mathbb{Q}_{p})$ (which is a subgroup of the automorphism group of a regular tree) into $\mathrm{Isom}(\mathbf{H}^{\infty})$.
More recently, simply connected non-positively curved symmetric spaces of infinite dimension and finite rank have been widely studied by Duchesne [15, 16, 17, 19]; his study includes the classification ([17], Theorem 1.8), the determination of their isometry groups ([19], Theorem 3.3), and a result about continuous representations of higher rank Lie groups into these isometry groups ([19], Theorem 1.1), stating that every such representation, provided it has no totally isotropic invariant subspace, splits as a direct sum of finite-dimensional representations and a unitary representation.
In another paper [44], we have studied representations of tree automorphism groups into Pontryagin spaces, and we have proved that for a wide class of tree automorphism groups, there are no continuous irreducible representations to Pontryagin spaces of index $\geq 2$, meaning that these groups admit no irreducible action on any simply connected non-positively curved irreducible symmetric space of infinite dimension and finite rank $\geq 2$.
Turning back to our theorem, it is convenient to define a family of possible target spaces $X$ that includes both finite- and infinite-dimensional spaces. To this end, we define the family $\mathcal{X}$ as follows:
@@
In the case where the space $X$ is finite-dimensional, the problem is already well understood. Indeed, let $f:G\rightarrow\mathrm{Isom}(X)$ be any continuous homomorphism. Since $\mathrm{Isom}(X)$ is a Lie group it has no small subgroups, and since $G$ is totally disconnected, by Van Dantzig’s theorem there is a basis of neighborhoods of the identity that are compact open subgroups. This implies that there exists a compact open subgroup of $G$ whose image in $\mathrm{Isom}(X)$ is trivial. As $G$ is almost simple, it follows from Theorem I.1.5.6 in [32] (main theorem in [43]) and Theorem I.2.3.1 in [32] (Proposition 6.14 in [8]) that every normal subgroup of $G$ is finite or cocompact. As $\ker{f}$ is a normal subgroup of $G$ containing a compact open subgroup, it has to be cocompact, meaning that $f$ has compact image. Since $X$ is a CAT(0) space, this implies the existence of a fixed point ([9], II.2.8).
We can start our proof of Theorem 1.1 with the reduction:
Proposition 1.5.
If we can prove Theorem 1.1 for $G=\mathrm{SL}_{3}(\mathbb{F})$ and $G=\mathrm{Sp}_{4}(\mathbb{F})$, then the general case follows.
Proof.
It is shown in [13] (Theorem 1.4, Proposition 1.2) that if $G$ is a simple linear algebraic group over a local field, then every action by isometries of $G$ on any metric space $X$ is either bounded, which means that the orbits are bounded, or metrically proper, which means that for every bounded subset $B\subset X$ the set $\{g\in G:gB\cap B\neq\emptyset\}$ has compact closure. The same holds for almost simple groups ([13], Lemma 3.1). In case the action is bounded, we can already deduce that it has a fixed point in $X$. In case it is metrically proper, we take a closed subgroup $H<G$ isomorphic to a quotient of $\mathrm{SL}_{3}(\mathbb{F})$ or $\mathrm{Sp}_{4}(\mathbb{F})$ by a finite subgroup. We can extend the action of $H$ on $X$ trivially to $\mathrm{SL}_{3}(\mathbb{F})$ or $\mathrm{Sp}_{4}(\mathbb{F})$; that action will have a fixed point $x\in X$. Now, the fact that the set $\{g\in G:gx=x\}$ has compact closure provides a contradiction since it contains $H$ which is closed and non-compact.
Proposition 1.5 allows, in the same way, to restrict Theorem 1.2 to the case $G=\mathrm{SL}_{3}(\mathbb{F})$.
It is enough to show that the action of a cocompact subgroup $\Gamma$ of $G=\mathrm{SL}_{3}(\mathbb{F})$ or $G=\mathrm{Sp}_{4}(\mathbb{F})$ is bounded. Our strategy will be to show that for a torsion-free cocompact lattice $\Gamma<G$, that is, a discrete torsion-free cocompact subgroup of $G$. As we have said, the existence of a cocompact lattice $\Gamma^{\prime}<G$ is guaranteed in our cases, and due to ([24], Theorem 2.7), there exists $\Gamma\lhd\Gamma^{\prime}$ of finite index that is also torsion-free. Moreover, as explained in [24], $\Gamma^{\prime}$ and hence $\Gamma$ is finitely generated.
Therefore, in order to prove Theorem 1.1 and Theorem 1.2, it is enough to show the following:
Theorem 1.6.
Let $\mathbb{F}$ be a non-archimedean local field. Let $G=\mathrm{SL}_{3}(\mathbb{F})$ or $G=\mathrm{Sp}_{4}(\mathbb{F})$; in the latter case assume that $\mathbb{F}$ has characteristic zero and its residue field $k$ has at least three elements. Let $\Gamma$ be a torsion-free cocompact lattice of $G$. Let $X\in\mathcal{X}$, as defined in Definition 1.3. Then any continuous action of $\Gamma$ by isometries on $X$ fixes a point in $X$.
We now present a short outline of the paper. First, we will show (Section 2) that every continuous action of $\Gamma$ by isometries on $X\in\mathcal{X}$ has a fixed point in $\overline{X}=X\cup\partial X$, where $\partial X$ is the boundary at infinity of $X$. This is done by generalizing a result that was proved by Wang in the finite-dimensional case [45], using a notion of harmonic functions from the set of vertices in the Bruhat-Tits building of $G$ to $X$. After this, we will show (Section 3) that if a locally compact group $\Gamma$ with Kazhdan’s Property (T) has the property that whenever it acts continuously by isometries on a space $X\in\mathcal{X}$ it fixes a point in $\overline{X}$, then the group actually has to fix a point in $X$. This will allow us to conclude our proof.
@@
In [45], Wang introduced a notion of admissible weight on simplicial complexes and proved that, if $\Gamma$ is the fundamental group of a finite simplicial complex with an admissible weight satisfying an eigenvalue condition, and $X$ is a complete simply connected finite-dimensional Riemannian manifold of nonpositive sectional curvature (as are the finite-dimensional spaces in our family $\mathcal{X}$), every continuous isometric action of $\Gamma$ on $X$ fixes a point in $\overline{X}$ ([45], Theorem 1.1). The list of eligible groups includes cocompact lattices in simply connected simple higher rank algebraic groups over non-archimedean local fields, provided that the cardinality of the residue field is large enough, with a bound that depends on the rank of the group ([45], Theorem 4.3). The results in our paper allow to show that, for groups $\Gamma$ that satisfy the assumptions of [45], every continuous isometric action on a space $X\in\mathcal{X}$ (including infinite-dimensional ones) has a fixed point in $\overline{X}$. In fact, one can follow the proof of our Theorem 2.1, showing existence of harmonic functions as we do in Section 2.2, and then show that harmonic functions are constant as in Section 3 of [45]. Furthermore, if the group $\Gamma$ has Kazhdan’s Property (T), as it is the case for lattices in connected almost simple higher rank algebraic groups over local fields ([4], Theorem 1.6.1, Theorem 1.7.1), then one can use our Theorem 3.4 to show that $\Gamma$ has to fix a point in $X$. This allows to extend our Theorem 1.1 to cocompact lattices in $G$ if the cardinality of the residue field is large enough, with a bound that depends on $\mathrm{rank}_{\mathbb{F}}(G)$.
@@
This paper is part of my Ph.D. project under the supervision of Prof. Nicolas Monod at EPFL (Lausanne, Switzerland). I am grateful to Nicolas for proposing this project, for giving me valuable advice and feedback, and for sharing with me deep insights of the theory. I am also grateful to Bruno Duchesne, Pierre Py, and the reviewers for their useful comments and suggestions on this paper, and in particular for the suggestion to generalize the original version of the theorem.
@@ 2 Fixing a point in $\overline{X}$
Let $\mathbb{F}$ be a non-archimedean local field. Let $G=\mathrm{SL}_{3}(\mathbb{F})$ or $G=\mathrm{Sp}_{4}(\mathbb{F})$; in the latter case assume that $\mathbb{F}$ has characteristic zero and its residue field $k$ has at least three elements. Let $\Gamma$ be a torsion-free cocompact lattice of $G$. Let $X\in\mathcal{X}$. Then every continuous action of $\Gamma$ by isometries on $X$ fixes a point in $\overline{X}$.
This result was proved by Wang in [45] in the case where $X$ is a finite-dimensional Riemannian manifold of non-positive sectional curvature, with an assumption that turns out to be more restrictive on the residue field. The proof of our Theorem 2.1 will closely follow the proof in [45], with some adaptations: most importantly, we need to take into account the fact that our space $X\in\mathcal{X}$ might be infinite-dimensional.
@@ 2.1 The building of $G$ and the eigenvalue property
First, we introduce the Bruhat-Tits building $\Sigma$ of the group $G$. Bruhat-Tits buildings were first introduced in [10]; more modern and readable treatments can be found in the books [1, 25]. In particular, we refer to Chapter 6.9 of [1] and Chapter 19 of [25] for the building of $\mathrm{SL}_{3}(\mathbb{F})$, and to Chapter 20.1 of [25] for the building of $\mathrm{Sp}_{4}(\mathbb{F})$. We also refer to the more recent papers [42, 41, 21] for a more detailed treatment of the building of $\mathrm{Sp}_{4}(\mathbb{F})$.
The Bruhat-Tits building $\Sigma$ of $G=\mathrm{SL}_{3}(\mathbb{F})$ is a Euclidean building of dimension two on which $G$ acts by isometries. It is a union of infinitely many Euclidean planes (apartments), each of them being tessellated by equilateral triangles (chambers). The action of $G$ is transitive on the set of pairs $(\mathcal{A},\mathcal{C})$ where $\mathcal{A}$ is an apartment and $\mathcal{C}$ is a chamber contained in $\mathcal{A}$. The situation is the same in the case $G=\mathrm{Sp}_{4}(\mathbb{F})$, except that the chambers are right isosceles triangles instead of equilateral. The vertices are divided into two categories: special and non-special, as described in [42, 41, 21]. Each chamber has two special vertices (at the $45^{\circ}$ angles) and one non-special vertex (at the $90^{\circ}$ angle).
@@
We describe the structure of the graph $Lk(v)$. In the case $G=\mathrm{SL}_{3}(\mathbb{F})$, as mentioned in [1] (6.9), it is the flag complex of $\mathrm{SL}_{3}(k)$, where $k$ is a finite field: the residue field of $\mathbb{F}$. This means that the vertices are points and lines in the projective plane over $k$, and whenever a point is contained in a line, the corresponding vertices are connected by an edge. If $q$ is the cardinality of the field $k$ (a power of a prime number), there are $2(q^{2}+q+1)$ vertices in total: $q^{2}+q+1$ points of $\mathbb{P}^{2}(k)$ and $q^{2}+q+1$ lines. Every vertex has degree $q+1$. Using the language of generalized polygons [22], the graph is a generalized triangle with $q+1$ lines through each point and $q+1$ points on each line.
In the case $G=\mathrm{Sp}_{4}(\mathbb{F})$, an analysis of the neighboring vertices of a vertex $v$ in the building was carried out in [21] (2.5). If $v$ is special, $Lk(v)$ is isomorphic to the flag complex of $\mathrm{Sp}_{4}(k)$, where $k$ is the residue field of $\mathbb{F}$. The vertices are points and isotropic lines in $\mathbb{P}^{3}(k)$, and there are $2(q^{3}+q^{2}+q+1)$ vertices in total: $q^{3}+q^{2}+q+1$ points and $q^{3}+q^{2}+q+1$ isotropic lines. Every vertex has degree $q+1$, and the graph is a generalized quadrangle with $q+1$ lines through each point and $q+1$ points on each line.
If $v$ is a non-special vertex in the building of $G=\mathrm{Sp}_{4}(\mathbb{F})$, the graph $Lk(v)$ is a complete bipartite graph with $q+1$ vertices on each side, or a generalized $2$-gon with $q+1$ points and $q+1$ lines.
We want to show a property of the graph $Lk(v)$ that corresponds to the first eigenvalue property in [45]. Let $v\in\Sigma(0)$ and let $Lk(v)$ be its link graph. Let $N$ be the number of vertices and $M$ be the number of edges in $Lk(v)$. After choosing an orientation for every edge in $Lk(v)(1)$, for every $f:Lk(v)(0)\rightarrow\mathbb{R}$ we can define $df:Lk(v)(1)\rightarrow\mathbb{R}$ by df((u,w))=f(u)f(w)df((u,w))=f(u)-f(w). We let $f||^{2}=\sum_{u\in Lk(v)(0)}|f(u)|^{2}$ and $df||^{2}=\sum_{(u,w)\in Lk(v)(1)}|f((u,w))|^{2}$. We want to establish a lower bound for $df||^{2}$ in terms of $f||^{2}$, for functions that satisfy $\sum_{u\in Lk(v)(0)}{f(u)}=0$.
Let $A$ be the adjacency matrix of $Lk(v)$. More explicitly, it is a matrix with $M$ rows and $N$ columns where, after an ordering of the vertices and of the edges has been chosen, the coefficient $A_{ij}$ is $1$ if the $i$-th edge starts from the $j$-th vertex, 1-1 if the $i$-th edge terminates at the $j$-th vertex, and 0 in all other cases. Then we can consider the function $f$ as a vector in $\mathbb{R}^{N}$ and the function $df$ as a vector in $\mathbb{R}^{M}$, so that we have $df=Af$. Therefore, we have
$df||^{2}=\langle df,df\rangle=\langle Af,Af\rangle=\langle f,A^{T}Af\rangle,$
@@
[…]
The calculation of minimal polynomials of the incidence matrices of generalized polygons has been carried out by Feit and Higman [22]. We also refer to [24], Proposition 7.10, for the complete list of eigenvalues in an equivalent setting. In our setting, the smallest non-zero eigenvalue of the matrix $B$ is:
  • $q+1-\sqrt{q}$, if $v$ is a vertex in the building of $\mathrm{SL}_{3}(\mathbb{F})$;
  • $q+1-\sqrt{2q}$, if $v$ is a special vertex in the building of $\mathrm{Sp}_{4}(\mathbb{F})$;
  • $q+1$, if $v$ is a non-special vertex in the building of $\mathrm{Sp}_{4}(\mathbb{F})$.
We may call $\lambda(v)$ the smallest non-zero eigenvalue of the matrix $B$ built from the link graph of the vertex $v$, as in the above list. As it is easy to see that the eigenspace corresponding to the eigenvalue 0 is the space of constant functions on $Lk(v)(0)$, the following proposition holds:
Proposition 2.2.
[…]
@@
The function $f$ in Proposition 2.2 takes values in $\mathbb{R}$, but the result also holds when it takes values in a general Hilbert space. To see this, it suffices to restrict to the finite-dimensional subspace spanned by the (finitely many) values of $f$, then apply the lemma component-wise.